The shape of Colorado is nearly a rectangle. The length is 100 miles than the width. The perimeter is about 1320 mi. Find the length and width of colorado.
thanks

Answer by SHUgrad05(58)

Perimeter of a rectangle: P=2l+2w
w=width of rectangle
l=length of rectangle; l=w+100
2*(w+100)+ 2w=1320 mi
2w+200+2w=1320mi
4w=1120 mi
w=280 mi
l=280+100=380 mi